OpenStudy (anonymous):

2. A car accelerates uniformly from +10.0 m/s to +40.0 m/s over a distance of 125 m. How long did it take to go that distance? Show all your work, including the equation used, given and unknown quantities, and any algebra required. Make sure your answer has the correct number of significant figures.

OpenStudy (anonymous):

distance (m) = change in speed (m/s) * change in time (s) 125 m = (40-10 m/s) * (seconds it took for the distance) 125 / 30 = time in seconds

OpenStudy (anonymous):

given: overall distance = 125 m final speed = +40 m/s initial speed = +10 m/s time = unknown s

OpenStudy (anonymous):

and that is all? Don't I have to calculate for acceleration too?

OpenStudy (anonymous):

Acceleration's not needed. For what's given, all you need is that equation to find the time, at least based on equations. Unless... I'm wrong. Which may likely be. \[v^{2} = vi^{2} + 2a (\Delta x)\] where: v = vfinal = +40m/s vi = vinitial = +10m/s using the equation, you can find acceleration \[v = vi + at\] using equation with v, vi, & a, you can find t

OpenStudy (anonymous):

Got it! Thank You :)